OpenStudy (anonymous):

find \[ {dy \over dx}xsin2x\]

OpenStudy (anonymous):

@.Sam.

sam (.sam.):

Differentiate xsin2x?

OpenStudy (anonymous):

use chain rule

OpenStudy (anonymous):

yeh

OpenStudy (cwrw238):

you also need the product rule f(uv) = uv' + vu' let u = x and v = sin2x

OpenStudy (anonymous):

(uv)'=u'v+uv' here u=x and v=sin2x

sam (.sam.):

dy/dx=x(cos(2x)(2))+(sin2x)

OpenStudy (diyadiya):

\[x \ \frac{d}{dx}\sin(2x)+ \sin(2x) \ \frac{d}{dx}x\]

OpenStudy (diyadiya):

\[(x )\ 2 \cos(2x)+\sin(2x)\]

OpenStudy (anonymous):

@Diyadiya I thought we should n't give away answers

OpenStudy (diyadiya):

huh ? @vamgadu not answers ..*Final answer